# Enhancing Transport Barriers with Swimming Microorganisms in Chaotic Flows

### v1

@@ 3.2 Numerical Simulations
To understand how swimming microorganisms interact with an imposed time-periodic flow, we perform numerical simulations of swimming particles using the experimentally measured velocity field. Microorganisms are modelled as axisymmetric ellipsoids with a constant swimming speed $v_{s}$, in the direction $\mathbf{q}$ along their symmetry axis. The swimmer’s position $\mathbf{x}$ is modelled as:
@@
where $\mathbf{v}_{f}$ is experimentally measured fluid velocity in the active suspension, and $v_{s}=20~{}\umu$m/s is the swimming speed of E. coli. As a control, simulations of elongated passive particles are also performed with $v_{s}=0$ in the velocity field measured in the buffer. The swimmer’s orientation is described by the Jeffery’s equation (Jeffery & Filon, 1922):
$\dot{\mathbf{q}}=[\mathbf{\Omega}(\mathbf{x},t)+\Lambda\mathbf{D}(\mathbf{x},t)]\mathbf{q}-\Lambda[\mathbf{q}\cdot\mathbf{D}(\mathbf{x},t)\mathbf{q}]\mathbf{q},$ (5)
@@
where $\mathbf{D}$ and $\mathbf{\Omega}$ are the symmetric and skew-symmetric parts of the velocity gradient tensor, $\nabla\mathbf{v}_{f}$. Here, $\Lambda=(1-\alpha^{2})/(1+\alpha^{2})$ is a shape factor, with $\alpha$ being the swimmer aspect ratio; the factor is roughly $\Lambda\approx 0.88$ for E. coli.
@@
Initially, both passive and active particles are uniformly distributed in the flow field with random orientations. As the simulations begin, passive and active particles start to develop complex patterns following the morphology of Lagrangian vortices (see Movie 3). Figures 3(a) and 3(b) show the spatial distribution of passive and active particles at $N=150$, respectively. The color code in the plots represents the local particle number density, $\rho_{N}$, normalized by the initial number density, $\rho_{0}$. We find that active particles are mostly depleted within the LVBs (green contours), suggesting that Lagrangian vortices repel elongated swimmers. This is in contrast to the passive case in which LVBs trap passive particles and lead to particle accumulation within the Lagrangian vortices. This results suggests the repulsion of active particles by the Lagrangian vortices is not a generic feature of the velocity field itself. Rather, it is the result of the interaction between Lagrangian vortices and elongated swimmers.
@@
The repulsion of elongated swimmers by the LVBs is quantified by the radial distribution function: $g(r)=\langle\rho_{N}(r)\rangle/\rho_{0}$, where the angle bracket denotes a radial average in all directions calculated from the elliptic fixed points. The function $g(r)$ represents the likelihood of finding a particle at a radial distance $r$ from the elliptic fixed point of a Lagrangian vortex. Figure 3(c) shows $g(r)$ for both passive and active particles in the Lagrangian vortex with the label “4” in Fig. 2. The dashed line in Fig. 3(c) is a nominal radius of Lagrangian vortex, defined as: $r_{n}=\sqrt{A_{S}/\pi}$, where $A_{S}$ is the area enclosed by the LVBs. We find that the likelihood of finding an active particle within the nominal radius of a Lagrangian vortex is smaller than that of a passive particle, and vice versa for outside the nominal radius. This result further corroborates the repulsion of elongated active particles by Lagrangian vortices.
@@
Figure 3: Spatial distributions of (a) passive particles, and (b) active particles, at time $N=150$. Particles are colored by their normalized local number density, $\rho_{N}/\rho_{0}$. Passive particles accumulate inside and active particles accumulate outside the LVBs (green contours). (c) Radial distribution function $g(r)$ calculated from the elliptic fixed points of the Lagrangian vortex with the label “4”; the dashed line is the nominal radius of the vortex. (d) Spatially averaged normalized number density within the LVBs, $\langle\rho_{N}\rangle_{S}/\rho_{0}$, as a function of time $N$ for the same vortex as in (c).
@@
To quantify the time evolution of the repulsion of elongated swimmers, we calculate the average number density within the LVBs, $\langle\rho_{N}\rangle_{S}$, normalized by the initial number density $\rho_{0}$. Figure 3(d) shows $\langle\rho_{N}\rangle_{S}/\rho_{0}$ as a function of time $N$. Results show that $\langle\rho_{N}\rangle_{S}$ decreases non-monotonically with time for active particles, and drops to only 20% of the initial number density $\rho_{0}$ at $N=200$. This suggests that most active particles escape and accumulate outside the Lagrangian vortex as mixing progresses. In the control case, $\langle\rho_{N}\rangle_{S}$ for passive particles increases with time and reaches four times of the initial number density at $N=200$, which suggests the repulsion of elongated particles by LVBs is not an intrinsic feature of the velocity field. Although it has been observed that symmetric and circular Eulerian vortices can repel elongated swimmers (Torney & Neufeld, 2007; Ran et al., 2021; Qin & Arratia, 2022), here, we extend these results in a more general case of asymmetric and non-circular vortices from Lagrangian criteria.
@@
Figure 4: (a) Stroboscopic trajectories of passive (blue) and active (red) particles that are initially inside the Lagrangian vortices illustrated by the TRA field. While sharing the same initial condition, passive particles remain trapped in the vortices and active particles spiral outward and escape. The trajectories of active particles are colored by their time, which appear more reddish as time progresses. (b) The PDFs of the inner product of the particle orientation vector, $\mathbf{q}$, and the tangent vector of members of elliptic LCSs, $\mathbf{t}$, as defined in the Inset. The initial condition ($N=0$) is approximately a uniform distribution for particles with random initial orientations. The PDFs at a later time ($N=50$) show that both passive and active particles preferentially align with members of elliptic LCSs at $\mathbf{q}\cdot\mathbf{t}=\pm 1$.
@@
To gain further insights into how swimming particles are repelled by Lagrangian vortices, we plot the stroboscopic trajectories and orientations of passive and active particles as a function of time, as shown in Fig. 4(a). The Lagrangian vortices are illustrated by the TRA field (color map); passive and active particles are shown as blue and red bars, respectively. The trajectories of passive particles are blue, while active particles’ trajectories are colored by elapsed time that become more reddish with time. Despite sharing the same initial conditions for position and orientation, passive particles remain trapped in the Lagrangian vortices while active particles spiral outward with time and escape the vortices. We note that the swimming number of the active particles, $\Phi=v_{s}/U$, is at the order of $10^{-2}$ in the simulations, which suggests that the swimming speed is fairly negligible compared to the flow speed. However, even such (relatively) low swimming speed can drive active particles out of equilibrium and escape the Lagrangian vortices. That is, even small levels of swimming activity is enough to initiate the expulsion process. In addition, we find that the orientations of both passive and active particles tend to align with the level curves of the TRA field. Since the convex TRA contours locate the nested families of elliptic LCSs, this result indicates elongated particles—whether self-propelled or not—preferentially align with the members of elliptic LCSs. The difference is that the self-propulsion of active particles causes them to move outward and leave the vortices in a spiral manner.
@@
Next, we test the alignment of the elongated particles, we calculate the inner product of the particle orientation vector, $\mathbf{q}$, and the tangent vector of the elliptic LCSs, $\mathbf{t}$. The inset of Fig. 4(b) shows a schematic of the two vectors. Since elliptic LCSs are nested families, we partition the elliptic LCSs into 10 members using the convex TRA contours. The inner product is calculated between the orientation vector of an elongated particle and the tangent of the nearest member of elliptic LCSs. Figure 4(b) shows probability density functions (PDF) of $\mathbf{q}\cdot\mathbf{t}$ for both passive and active particles at two different times. At $N=0$, the initial condition of the PDFs shared by both passive and active particles is approximately a uniform distribution due to the random initial orientations. At a later time of $N=50$, the PDFs become bimodal at $\mathbf{q}\cdot\mathbf{t}=\pm 1$, suggesting a parallel or anti-parallel alignment between $\mathbf{q}$ and $\mathbf{t}$; we do not find a substantial difference between the PDFs of passive and active particles, indicating self-propulsion does not play a significant role in affecting the alignment of elongated particles with elliptic LCSs. However, self-propulsion causes active particles to travel toward the outer members of the elliptic LCSs in both parallel and anti-parallel alignment configurations (see Movie 4), resulting in the spiral escaping trajectories of the active particles in Fig. 4(a). Consequently, the interplay between self-propulsion and the alignment of elongated particles with elliptic LCSs leads to the depletion of active particles inside the elliptic LCSs, that is, the accumulation of active particle outside the Lagrangian vortex boundaries (LVBs). This accumulation of microswimmers outside the LVBs can further obstruct dye transport into the Lagrangian vortices, which is responsible for the observed transport barriers.
@@ 4 Conclusion
In this manuscript, we investigate the interaction between swimming microorganisms and Lagrangian coherent vortices known as the elliptic LCSs in time-periodic flows in experiments and in simulations. Our results show that even small amounts of swimming activity can affect (i) the dynamics of active particles in the flow (Fig. 4) and consequently (ii) the mixing and transport of passive scalars (Fig 1) in chaotic flows. Experiments show that the interaction between organisms and elliptic LCSs leads to transport barriers through which the tracer flux is significantly reduced. Using the Poincaré map and the TRA field, we show that these transport barriers coincide with outermost member of elliptic LCSs, or Lagrangian vortex boundaries (LVBs). To further understand the formation of the transport barriers, we perform numerical simulations of elongated microswimmers in experimentally measured velocity fields. Results show that elliptic LCSs can repel elongated swimmer and lead to swimmer accumulation outside LVBs. This accumulation of microswimmers effectively reduces the transport into elliptic LCSs. We further show that the interplay between self-propulsion and the preferential alignment of elongated particle with the tangents of elliptic LCSs leads swimmers to escape the Lagrangian vortices. Overall, our results allow quantitative prediction of the Lagrangian transport of microorganisms and passive tracer quantities (e.g., temperature, oxygen, and nutrients) in chaotic flows with nontrivial vortex structures. Although there have been previous studies on the interaction between microorganisms and LCSs (Khurana & Ouellette, 2012; Dehkharghani et al., 2019; Ran et al., 2021; Qin & Arratia, 2022; Si & Fang, 2021, 2022; Yoest et al., 2022), our work extend study of swimmer-LCS interaction to elliptic LCSs (i.e., vortex-like dynamical structures). From a practical perspective, our results may be useful in understanding the organic matter flux in the oceans, algal blooms in lakes, and harmful bacterial infections.

### v2

@@ 3.2 Numerical Simulations
To understand how swimming microorganisms interact with an imposed time-periodic flow, we perform numerical simulations of swimming particles using the experimentally measured velocity field. Microorganisms are modeled as axisymmetric ellipsoids with a constant swimming speed $v_{s}$, in the direction $\mathbf{q}$ along their symmetry axis. The swimmer’s position $\mathbf{x}$ is modelled as:
@@
where $\mathbf{v}_{f}$ is experimentally measured fluid velocity in the active suspension, and $v_{s}=20~{}\umu$m/s is the swimming speed of E. coli. As a control, simulations of elongated passive particles are also performed with $v_{s}=0$ in the velocity field measured in the buffer. The swimmer’s orientation is described by Jeffery’s equation (Jeffery & Filon, 1922):
$\dot{\mathbf{q}}=[\mathbf{W}(\mathbf{x},t)+\Lambda\mathbf{D}(\mathbf{x},t)]\mathbf{q}-\Lambda[\mathbf{q}\cdot\mathbf{D}(\mathbf{x},t)\mathbf{q}]\mathbf{q},$ (5)
@@
where $\mathbf{D}$ and $\mathbf{W}$ are the symmetric and skew-symmetric parts of the velocity gradient tensor, $\nabla\mathbf{v}_{f}$. Here, $\Lambda=(1-\alpha^{2})/(1+\alpha^{2})$ is a shape factor, with $\alpha$ being the swimmer’s aspect ratio. We assume $\alpha=0.25$ for rod-shaped E. coli, which leads to $\Lambda\approx 0.88$.
@@
Initially, both passive and active particles are uniformly distributed in the flow field with random orientations. As simulations begin, passive and active particles begin to develop complex patterns following the morphology of Lagrangian vortices (Movie S3). Figures 3(a) and 3(b) show the spatial distribution of passive and active particles at $N=150$, respectively. The color code in the plots represents the local particle number density, $\rho_{N}$, normalized by the initial number density, $\rho_{0}$. We find that active particles deplete within and aggregate outside the LVBs. This suggest that Lagrangian vortices repel elongated swimmers. By contrast, no depletion is observed for passive particles. Accumulation is not expected for passive particles in a 2D incompressible flow, and aggregation of non-motile bacteria has not been observed in previous experiments (Ran et al., 2021). The accumulation of passive particles seen here is likely due to a small departure from an (ideal) divergence-free experimental velocity field (see Supplementary Materials). The passive particle simulation serves as a control to show that the repulsion of active particles by the Lagrangian vortices is not an intrinsic feature of the experimental velocity field itself. Rather, it is the result of the interaction between Lagrangian vortices and elongated swimmers.
@@
The repulsion of rod-shaped swimmers by the LVBs is quantified by the radial distribution function: $g(r)=\langle\rho_{N}(r)\rangle/\rho_{0}$, where the angle bracket denotes a radial average in all directions calculated from the elliptic fixed points. The function $g(r)$ represents the probability of finding a particle at a radial distance $r$ from the elliptic fixed point of a Lagrangian vortex. Figure 3(c) shows $g(r)$ for both passive and active particles in the Lagrangian vortex with the label “4” in Fig. 2. The dashed line in Fig. 3(c) is a nominal radius of Lagrangian vortex, defined as: $r_{n}=\sqrt{A_{S}/\pi}$, where $A_{S}$ is the area enclosed by the LVBs. We find that the likelihood of finding an active particle within the nominal radius of a Lagrangian vortex is smaller than that of a passive particle and vice versa for outside the nominal radius. This result further corroborates the repulsion of elongated active particles by Lagrangian vortices.
@@
Figure 3: Spatial distributions of (a) passive particles, and (b) active particles, at time $N=150$. Particles are colored by their normalized local number density, $\rho_{N}/\rho_{0}$. Active particles are depleted from the vortex and accumulate outside the LVBs (green contours), while this depletion is not present for passive particles. (c) Radial distribution function $g(r)$ calculated from the elliptic fixed points of the Lagrangian vortex with the label “4”; the dashed line is the nominal radius of the vortex. (d) Spatially averaged normalized number density within the LVBs, $\langle\rho_{N}\rangle_{S}/\rho_{0}$, as a function of time $N$ for the same vortex as in (c).
@@
To quantify the time evolution of the repulsion of elongated swimmers, we calculate the average number density within the LVBs, $\langle\rho_{N}\rangle_{S}$, normalized by the initial number density $\rho_{0}$. Figure 3(d) shows $\langle\rho_{N}\rangle_{S}/\rho_{0}$ as a function of time $N$. Results show that $\langle\rho_{N}\rangle_{S}$ decreases nonmonotonically with time for active particles, and drops to only 20% of the initial number density $\rho_{0}$ at $N=200$. This suggests that most active particles escape and accumulate outside the Lagrangian vortex as mixing progresses. In the control case, $\langle\rho_{N}\rangle_{S}$ for passive particles increases with time, which again suggests the repulsion of elongated particles by LVBs is not an intrinsic feature of the velocity field itself. Although symmetric and circular Eulerian vortices have been observed to repel elongated swimmers (Torney & Neufeld, 2007; Ran et al., 2021; Qin & Arratia, 2022), here we extend these results to a more general case of asymmetric and non-circular vortices from Lagrangian criteria.
@@
Figure 4: (a) Stroboscopic trajectories of passive (blue) and active (red) particles that are initially inside the Lagrangian vortices illustrated by the TRA field. While sharing the same initial condition, passive particles remain trapped in the vortices and active particles spiral outward and escape. The trajectories of active particles are colored by their normalized time (or numbers of periods), $N/N_{\mathrm{tot}}$, where $N_{\mathrm{tot}}$ is the total time duration of the trajectories. (b) The PDFs of the inner product of the particle orientation vector, $\mathbf{q}$, and the tangent vector of the elliptic LCSs in the direction of the vortex circulation, $\mathbf{t}$, as defined in the inset. The initial condition ($N=0$) is approximately a uniform distribution for particles with random initial orientations. The PDFs at a later time ($N=50$) show that both passive and active particles preferentially align with members of elliptic LCSs at $\mathbf{q}\cdot\mathbf{t}=\pm 1$. The PDF of active particles is slightly biased towards the positive peak of $\mathbf{q}\cdot\mathbf{t}=+1$. (c) Schematic of a “bacterial porous medium” formed by cells aligning and accumulating outside a Lagrangian vortex boundary (LVB). Dye transport is hindered as it diffuses through the porous media.
@@
To gain further insights into how swimming particles are repelled by Lagrangian vortices, we plot the stroboscopic trajectories and orientations of passive and active particles as a function of time, as shown in Fig. 4(a). The Lagrangian vortices are illustrated by the TRA field (color map); passive and active particles are shown as blue and red bars, respectively. The trajectories of the passive particles are blue, while the trajectories of the active particles are colored by their normalized time or number of periods, $N/N_{\mathrm{tot}}$, where $N_{\mathrm{tot}}$ is the total time duration of the trajectories. Despite sharing the same initial conditions for position and orientation, passive particles remain trapped in the Lagrangian vortices while active particles spiral outward with time and escape the vortices. We note that the swimming number of the active particles, $\Phi=v_{s}/U$, is of the order of $10^{-2}$ in the simulations, suggesting that the swimming speed is fairly negligible compared to the flow speed. However, even such a (relatively) low swimming speed can drive active particles out of equilibrium and escape the Lagrangian vortices. That is, even small levels of swimming activity is enough to initiate the expulsion process. In addition, we find that the orientations of both passive and active particles tend to align with the level curves of the TRA field. Because the convex TRA contours locate the nested families of elliptic LCSs, this result indicates elongated particles—whether self-propelled or not—preferentially align with the members of elliptic LCSs. The difference is that the self-propulsion of active particles causes them to move outward and leave the vortices in a spiral manner.
@@
Next, we test the alignment of the elongated particles by calculating the inner product of the particle orientation vector, $\mathbf{q}$, and the tangent vector of the elliptic LCSs in the direction of vortex circulation, $\mathbf{t}$. The inset of Fig. 4(b) shows a schematic of the two vectors. Since elliptic LCSs are nested families, we partition the elliptic LCSs into 10 members using the convex TRA contours. The inner product is calculated between the orientation vector of an elongated particle and the tangent of the nearest member of elliptic LCSs. Figure 4(b) shows probability density functions (PDFs) of $\mathbf{q}\cdot\mathbf{t}$ for both passive and active particles at two different times. At $N=0$, the initial condition of the PDFs shared by both passive and active particles is approximately a uniform distribution due to the random initial orientations. At a later time of $N=50$, the PDFs of both passive and active particles become bimodal at $\mathbf{q}\cdot\mathbf{t}=\pm 1$, suggesting a parallel or anti-parallel alignment between $\mathbf{q}$ and $\mathbf{t}$. We find that the PDF of active particles shows a slight bias towards the positive peak at $\mathbf{q}\cdot\mathbf{t}=+1$, while the bias is not present for the PDF of passive particles. This suggests that active particles prefer to swim parallel to the elliptic LCSs in the direction of vortex circulation rather than swimming against the circulation. The bias in parallel and anti-parallel alignments is also found for active particles of different swimming speeds (see Supplementary Materials). The self-propulsion of active particles can cause them to travel toward the outer members of the elliptic LCSs in both parallel and anti-parallel alignment configurations (see Movie S4), resulting in the spiral escaping trajectories of the active particles shown in Fig. 4(a). Consequently, the interplay between self-propulsion and the alignment of elongated particles with elliptic LCSs leads to the depletion of active particles inside the elliptic LCSs, that is, the accumulation of active particles outside the Lagrangian vortex boundaries (LVBs). This accumulation of microswimmers outside the LVBs can further obstruct dye transport into the Lagrangian vortices, which is responsible for the observed transport barriers.
3.3 Physical Mechanism
We now propose a potential mechanism to explain the observed reduction in dye transport into Lagrangian vortices. We posit that the alignment and accumulation of bacteria outside the LVBs form an effective “bacterial porous medium” that can impede the diffusion of dye molecules, as schematically shown in Fig. 4(c). In the absence of bacteria, dye should enter LVBs primarily by diffusion and marginally by inertial effect due to (low but) nonzero Reynolds number ($Re\sim 10^{1}$). Here, we will focus on diffusive transport around the outside boundary of the LVBs where bacteria accumulate [Fig. 4(c)]. The effective diffusivity of a molecular dye through a porous medium, $D_{\mathrm{eff}}$, can be estimated as follows (Grathwohl, 1998):
$D_{\mathrm{eff}}=\frac{D_{0}\epsilon\delta}{\tau},$ (6)
where $D_{0}$ is the intrinsic diffusivity of the dye molecules, $\epsilon$, $\delta$, and $\tau$ are the porosity, constrictivity, and tortuosity of the porous media, respectively. The porosity, $\epsilon$, is the volume fraction of the void spaces (that is, pores) in the medium relative to its total volume; $\epsilon$ ranges from 0 to 1. For our bacterial medium, the quantity $\epsilon$ sums up to unity with the bacterial volume fraction $\phi_{b}$ such that $\epsilon=1-\phi_{b}$. Bacterial accumulation around LVB leads to a local increase in $\phi_{b}$ of approximately tenfold (as shown by our simulations) such that $\phi_{\mathrm{LVB}}\approx 10\phi_{b}=0.05$. Thus, the porosity of the bacterial medium is $\epsilon=1-\phi_{\mathrm{LVB}}\approx 0.95$. This leads to only a 5% decrease in $D_{\mathrm{eff}}$, which cannot account for the observed reduction in dye transport. Similarly, the constrictivity $\delta$ is a dimensionless parameter ranging from 0 to 1, which captures the hindrance to which a diffusing substance is subjected to when travelling through narrow pores. It becomes important only if the size of the diffusing molecules is comparable to that of the pores (Stenzel et al., 2016; Bini et al., 2019). Here, the pore size, $\mathcal{O}(1~{}\umu$m), is much greater than the molecular size, $\mathcal{O}$(1 nm), and therefore $\delta\approx 1$. Since the values of $\delta$ and $\epsilon$ are close to unity, we do not expect them to play a significant role in the reduction of dye transport.
Next, we examine the role of tortuosity ($\tau$), which compares the (tortuous) pathway of molecular diffusion in a porous medium to its pathway in an unrestricted medium (Bini et al., 2019; da Silva et al., 2022). The quantity $\tau$ can be defined as (Grathwohl, 1998; Holzer et al., 2013; da Silva et al., 2022):
$\tau=\frac{\langle L_{s}\rangle}{L_{0}},$ (7)
where $L_{0}$ is the length of the straightest path and $\langle L_{s}\rangle$ is the ensemble average of all possible tortuous path of diffusion. Note that $\tau\geq 1$ since $\langle L_{s}\rangle\geq L_{0}$. We consider the transverse diffusion of dye molecules across a bacterial porous medium, as sketched in Fig. 4(c), where $x_{1}$ denotes the direction of the transverse diffusion and $x_{2}$ denotes the direction of cell body alignment with the tangent of the elliptic LCSs. During the diffusion process, each time the (dye) molecule encounters an obstacle (i.e., a bacterium), its path will be deflected by on average half the body length of a bacterium, $l_{E}/2$. Also, dye diffusing a distance $L_{0}$ (in the bacterial porous medium) will encounter bacterial cells on average $L_{0}\sigma_{N}$ times, where $\sigma_{N}$ is the cell number density. We can then estimate the average transverse and longitudinal path-lengths, $\langle L_{s}\rangle_{11}$ and $\langle L_{s}\rangle_{22}$, for the dye to diffuse across a distance $L_{0}$ in the $x_{1}$ and $x_{2}$ directions as:
$\langle L_{s}\rangle_{11}=L_{0}+L_{0}\sigma_{N}l_{E}/2,\ \langle L_{s}\rangle_{22}=L_{0}+L_{0}\sigma_{N}\alpha l_{E}/2,$ (8)
where $\alpha=d_{E}/l_{E}$ is the aspect ratio between the average cell diameter $d_{E}$ and cell length $l_{E}$. For rod-shaped bacteria such as E. coli, we notice that $\alpha<1$ and thus $\langle L_{s}\rangle_{22}<\langle L_{s}\rangle_{11}$. This suggests that diffusion is anisotropic in the bacterial porous medium. We can now express the tortuosity for the transverse and longitudinal diffusion as:
$\tau_{11}=\frac{\langle L_{s}\rangle_{11}}{L_{0}}=1+\sigma_{N}l_{E}/2,\ \tau_{22}=\frac{\langle L_{s}\rangle_{22}}{L_{0}}=1+\alpha\sigma_{N}l_{E}/2.$ (9)
In our experiments, the local bacterial volume fraction (around LVBs) is $\phi_{\mathrm{LVB}}\approx 0.05$, which corresponds to a local cell number density of $\rho_{N}\approx 1.25\times 10^{11}$ cells/mL, and a local cell number density of $\sigma_{N}\approx 5,000$ cells/cm. If we consider $l_{E}$ to be the bacterium body length ($\approx 2~{}\umu$m), then $\tau_{11}=1+\sigma_{N}l_{E}/2=1.5$. If instead we consider the ”total” bacterium length (cell body plus flagella), $L_{\mathrm{total}}\approx 7~{}\umu$m (Patteson et al., 2015), then $\tau_{11}=1+\sigma_{N}L_{\mathrm{total}}/2=2.75$. Using the above estimates, we expect $1.5\leq\tau_{11}\leq 2.75$. The transverse effective diffusivity can be estimated using Eq. 6, and this yields $0.346\leq D_{11}/D_{0}\leq 0.633$. Therefore, a dye diffusing through this bacterial porous media would experience a 33%–66% decrease in the transverse effective diffusivity, which qualitatively corresponds to the decrease in dye transport into the LVBs. Note that Eq. 9 estimates an anisotropic diffusivity, $D_{11}<D_{22}$, for rod-shaped bacteria ($\alpha<1$). Bacterial flagellar motion, in addition, can displace dye molecules further in the $x_{2}$ direction, which would lead to an increase in $D_{22}$. This would lead to a more complex behavior than the one described here. Nevertheless, the concept of anisotropic diffusion in a bacterial porous media seems to capture, at least qualitatively, the observed decrease in scalar transport into the Lagrangian vortices.
@@ 4 Conclusion
In this manuscript, we investigate the interaction between swimming microorganisms and Lagrangian coherent vortices known as the elliptic LCSs in time-periodic flows in experiments and in simulations. Our results show that even small amounts of swimming activity can affect (i) the dynamics of active particles in the flow (Fig. 4) and consequently (ii) the mixing and transport of passive scalars (Fig. 1) in chaotic flows. Experiments show that the interaction between organisms and elliptic LCSs leads to transport barriers through which the tracer flux is significantly reduced. Using the Poincaré map and the TRA field, we show that these transport barriers coincide with outermost member of elliptic LCSs, or Lagrangian vortex boundaries (LVBs). To further understand the formation of the transport barriers, we perform numerical simulations of elongated microswimmers in experimentally measured velocity fields. Results show that elliptic LCSs can repel elongated swimmer and lead to swimmer accumulation outside LVBs. This accumulation of microswimmers effectively reduces the transport into elliptic LCSs. We further show that the interplay between self-propulsion and the preferential alignment of elongated particle with the tangents of elliptic LCSs leads swimmers to escape the Lagrangian vortices. Overall, our results allow quantitative prediction of the Lagrangian transport of microorganisms and passive tracer quantities (e.g., temperature, oxygen, and nutrients) in chaotic flows with nontrivial vortex structures. Although there have been previous studies on the interaction between microorganisms and LCSs (Khurana & Ouellette, 2012; Dehkharghani et al., 2019; Ran et al., 2021; Qin & Arratia, 2022; Si & Fang, 2021, 2022; Yoest et al., 2022), our work extend study of microorganism-LCSs interaction to elliptic LCSs (i.e., vortex-like dynamical structures). From a practical perspective, our results may be useful in understanding the organic matter flux in the oceans, algal blooms in lakes, and harmful bacterial infections.